A candy mixture is created with 2 types of candy, one costing $4 per pound and the other $3.50 per pound. How much of each type is needed for a 5 pound box that coast $18?
Answer by jessica43(140)

To solve this problem you need to write two separate equations using what you know.
First, you know that there are two types of candy and together they must total 5 pounds:
A + B = 5
This can be rewritten as A = 5 - B.
Second, you know that candy A costs $4 per pound and candy B costs $3.50 per pound and together they must total $18:
4A + 3.50B = 18
Now you can plug in the rewritten first equation into the second equation and solve for B:
4A + 3.50B = 18
4(5 - B) + 3.50B = 18
20 - 4B + 3.50B = 18
20 - 0.5B = 18
-0.5B = -2
B = 4
So you need 4 pounds of candy B (at $3.50 per pound). Since you need 5 pounds total, this means you need only 1 pound of candy A (at $4 per pound).